OpenStudy (rane):

at what dstance from the earth, in a direct line, between the sun and the earth will the gravitational attraction between the 2 bodies be the same?

OpenStudy (rane):

@DoritosHero

OpenStudy (rane):

@Zale101

Parth (parthkohli):



Parth (parthkohli):

Let's say that the distance of the body from Earth is \(x\)

Parth (parthkohli):

What is the distance from the Earth to the Sun?

OpenStudy (rane):

there is no values given

Parth (parthkohli):

OK, let's assume 150,000,000,000 meters.

OpenStudy (rane):

k

Parth (parthkohli):

If the distance of the body to the Earth is \(x\), then what is the distance from the body to the Sun?

OpenStudy (rane):

150,000,000,000 meters - x

Parth (parthkohli):

You got it. Now, we can use the gravitational force formula\[F = \dfrac{GMm}{R^2}\]What are the masses of Earth and Sun given as?

Parth (parthkohli):

For the mass of Earth, let's say \(6 \times 10^{24} ~ \rm kg\). For the mass of Sun, \(2 \times 10^{30}~\rm kg\).

Parth (parthkohli):

Got it now?

Parth (parthkohli):

We just want equal gravitational force... so?